Currently browsing tag


TEM is down

June 3, 2014. The TEM vacuum is bad so I will regenerate the system overnight.  If all goes well it should be okay to use tomorrow.  I will send out an update then.

TEM status

November 7, 2013. Installation of the new(er) TEM is proceeding as we planned. We will have the system under vacuum where it will stay through next week. Final installation steps will resume Monday, Nov 18. We won’t remove any parts from the 4000 TEM until that week so, if you …


TEM down

September 3, 2013. There was a power outage in Houghton over the weekend that shut down the TEM. I restarted it this morning but the vacuum is slow to recover. I’ll send another message when it is back online.


TEM status – online

June 6, 2013. The TEM is repaired and, after the gun vacuum improves overnight, should be ready for the first appointment (Yang). Note that one of the standard sample holders was bent somehow. I’ll send it in for repair. Be careful – repairs like this lead to higher use charges.


FESEM is down

March 27, 2013. The FESEM is down with a failed vacuum pump. Hitachi si sending another that will arrive about lunchtime tomorrow. The new pump should repair the problem. I will keep you informed of progress.

Contaminated Sputter Coater

February 27, 2012. The sputter coating system is contaminated and the coatings are of poor quality.  Please don’t use it until I get it cleaned up.  I’ll post a follow-up message to let you know when it is ready. Why is the coater becoming contaminated?  It happens when wet samples …